HUBBLE SPACE TELESCOPE DAILY REPORT #967 PERIOD COVERED: 0800 EST Daily Status Report as of 251/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WFC/PC 5026 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor) The Wide Field Camera and the Planetary Camera were used to make two internal flat calibration images. These images will be used to look for measles and daisies (i.e., different forms of contamination), and to monitor scattered light. Both observations were executed as scheduled, and no problems were reported. The internal flat observations were reported to look normal. 1.2 Completed PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) The Planetary Camera was used to make two standard Earth calibrations images for flat fielding of PC observations. Medium and narrow filters were used, with and without neutral density filters. Both sets of observations were executed as scheduled. The first image was underexposed, and the second was reported to look normal. 1.3 Completed WFC/PC 5025 (WF/PC cycle 3 Calibration: Photometric Calibration: Monitor) The Wide Field Camera and planetary Camera were used to make observations of the ultraviolet star BD+75D325 using only single-chip readouts of W2 and P6. Six broad band filters were used. A total of twelve observations were taken. All twelve observations were reported to look normal. 1.4 Completed FOS/BL 3803 (Pluto's FUV spectrum: CO Abundance and FUV Survey) The Faint Object Spectrograph was used to attempt to make far ultraviolet observations of the planet Pluto to look for CO absorption in the Cameron Bands. This does not include its moon Charon. Unfortunately the baseline guide star acquisition failed, so the shutter never opened for the following observations. As a result, the planned SIP corrective maneuvers were canceled. Subsequent reacquisitions also failed. It was reported that the problem appeared to be a bad secondary guide star. No observations were taken. (See Operations Note 2.1) 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 2 Successful : 1 Scheduled Reacquisitions : 13 Successful : 4 FGS 1 & 3 failed acquisition of a guide star in Fine Lock at 249/1933z, and acquired successfully acquired Course Track. At 250/2010Z Acquisition completely failed, and course track was not obtained. An ARU/PRT was successfully uplinked to the spacecraft, but the FGSs could not located the guide stars. PMT counts were too low for observing. 2.2 FHST Updates: Scheduled : 5 Successful : 5 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete GHRS/2 3584 (Probing the Voids) 3.2 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.3 Complete WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) 3.4 Complete FOC/96 4153 (FOC UV Throughput Monitoring) 3.5 Complete FOS/BL 5053 (Cycle 2 FOS Location of Spectra: Y-Base Maps) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-